OpenStudy (anonymous):

Which of the following is an arithmetic sequence? A. –4, –2, 0, 2, … B. –8, –4, –2, –1, … C. –4, –3, –1, 2, … D. –4, 0, –4, 0, …

OpenStudy (eric_d):

Find the common difference,d

OpenStudy (eric_d):

Use \(T2-T1=T4-T3\)

OpenStudy (eric_d):

If the common difference of that particular option, then, it's an AP! @VicBrad5

OpenStudy (anonymous):

i know its not B im thinking that maybe its C

OpenStudy (eric_d):

are equal *

OpenStudy (eric_d):

Okay, will confirm it soon !

OpenStudy (eric_d):

Give me a sec !

OpenStudy (anonymous):

okay thank you

OpenStudy (eric_d):

It's A and D @VicBrad5

OpenStudy (anonymous):

wait how is it both?

OpenStudy (eric_d):

Let's get someone to confirm it ! @VicBrad5

OpenStudy (anonymous):

i feel as if it would be A more than D bc its an actual sequence

OpenStudy (eric_d):

@ribhu Pls confirm this. Tq.

OpenStudy (ribhu):

2a2 = a1+a2

OpenStudy (ribhu):

apply and check

OpenStudy (eric_d):

@VicBrad5

OpenStudy (anonymous):

It is A! :)

OpenStudy (ribhu):

yes @VicBrad5